OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is x – 7 and whose width is x + 12? Answer a.x2 – 84 b.x2 + 5x – 84 c.x2 – 84x + 5s d.x2 – 5x – 84

OpenStudy (anonymous):

area=l*w area=(x-7)(x+12)

OpenStudy (anonymous):

should be D

OpenStudy (anonymous):

foil it out...you get x^2 -12x + 7x -84 combine like terms you get x^2 -5x -84

OpenStudy (anonymous):

b.x2 + 5x – 84